OpenStudy (anonymous):

At a temperature of 280 K, the gas in a cylinder has a volume of 20.0 liters. If the volume of the gas is decreased to 10.0 liters, what must the temperature be for the gas pressure to remain constant? A. 140 K B. 273 K C. 560 K D. 5600 K

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

I am Sorry There was a mistake in my calculation..i just Found it nw PV = n R T P = (n R T1) /V1 P = (n R T2) / V2 comparing T1/V1 = T2/V2

OpenStudy (anonymous):

T2 = T1/V2 * V1

OpenStudy (anonymous):

So...That the Answer Would be 140K